OpenStudy (anonymous):

d/dx int_{1}^{sinx} sqrt{2+t^{4}}dt

OpenStudy (anonymous):

Use the chain rule. In this case the inner function is \(\sin x\)

OpenStudy (anonymous):

\[d/dx \int\limits_{1}^{sinx}\sqrt{2+t^{4}}dt\]

OpenStudy (anonymous):

sooo\[\sqrt{3}-\sqrt{2+\sin^{4}x}*cosx\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

What is with the \(\sqrt{3}\)?

OpenStudy (anonymous):

by F.T.C part II differentiate F(1)-F(sinx) correct? if so d/dx of F(1) = sqrt{2+{1^4}} = sqrt(3)?

OpenStudy (anonymous):

d/dx of F(1) = 0 because F(1) is constant.

OpenStudy (anonymous):

\(F(1)\) is a constant with respect to \(x\).

OpenStudy (anonymous):

that was my thoughts! thanks!